# Exploring the test of the no-hair theorem with LISA

### v1

@@ III.2 Templates
We consider two approaches where the recovery template takes different forms: the agnostic approach and the deviation approach. In the agnostic approach, we assume that the ringdown waveform is described by
@@
Here, complex frequencies, amplitudes, and phases can take any value. Note also that any dependence on the spheroidal harmonic is absorbed in the amplitude and the phase. In this way, no mode mixing is specified. For example, in this approach, it would not be possible to know how much of the $(2,2,0)$ QNM contribution comes from the spherical harmonic $(2,2)$ or the $(3,2)$. It differs from Eq. (1) as no assumption is made on the value of the complex frequency nor the spherical contribution to any QNM. Despite the fixed number of modes $k$, we can call this approach “agnostic”.
In the deviation approach, we assume the framework of GR as baseline but allow for a small “deviation” in the complex frequencies
@@ III.3 Bayesian analysis
where $\bm{\theta}$ is the vector of the source physical parameters, $M$ is the model or any other feature considered. In the numerator, we have the likelihood $\mathcal{L}(\theta)=p(d|\bm{\theta},M)$ and the prior of the parameters $\pi(\theta)=p(\bm{\theta}|M)$. In the numerator, $\mathcal{Z}=p(d|M)$ is the evidence, which is computed as the integral of the likelihood over the whole parameter’s hyper-volume. For a noise with a covariance matrix $\mathbf{C}$, the likelihood takes the form
@@
if the noise covariance is fixed. We drop the dependence on the model $M$ and we use the definition of the inner product in the time domain
@@
where $\rho(k)$, $k=|i-j|$, is the autocovariance function (ACF) that can be estimated from noise-only data in TD with a length longer than $N$ or as the inverse Fourier transform of the PSD. In our case, we use the latter option, generating the ACF from the LISA Science Requirements Document (SCiRD) [68] PSD
@@
Then, the inner product of Eq. (27) becomes a much faster product of vectors
@@
Figure 2: Posterior distribution of the pairs of complex frequencies in the spectrum map. Each mode ‘k’ is associated to one colour, purple, pink or green. The spectrum of the true BH is represented with coloured crosses with their QNM label nearby.
@@ IV.1 Agnostic approach
In the agnostic approach the parameters are $\bm{\theta}~{}=~{}\{A_{k},\phi_{k},\omega_{k},\tau_{k}\}$ with $k=1,2,3$ accounting for the three QNMs $[(2,2,0),(3,3,0),(4,4,0)]$. To avoid any degeneracy between the modes, we impose the condition of hyper-triangulation in the frequency. This condition restricts the second frequency to be larger than the first one, and the third to be larger than the second. Then, the uniform prior of each frequency decreases relative to the previous one, like an inverted triangle in the prior’s volume. The amplitudes have a logarithmic uniform prior in $[-23,\,-16]$, as well as the frequency in [5, 0]5 0[-5,\,0][ - 5 , 0 ] and the damping time in [0, 6]06[0,\,6][ 0 , 6 ]. The phase is allowed to take any value in the range [0, $[0,\,2\pi]$.
@@
In Fig. 1, we present the posterior distribution of the injection without noise in red, with noise in blue, and the injected values with black lines. Remember that the values of the deviations injected in the waveform have been introduced in Table 1. We show 12 parameters, 4 for each QNM. In general, the Gaussian distributions converge to the true values, with some minor fluctuations in the noisy case, as expected. In the figure we have already labeled the name of the QNMs “$k$” since we know them from the injection. However, we will not know which modes are present in the future LISA data analysis. Therefore, one must first find the QNM corresponding to each label $k=lmn$.
@@
Figure 4: Computed mass for spins in the range [0.8-0.9] with Eq.34 with a $99\%$ confidence level, from the estimated mean values of ($\omega_{lmn},\tau_{lmn}$)
One possible way to check for consistency between QNMs is to trace back the mass and spin of the remnant BH, using for example the fittings from [23]:
@@ IV.1.1 Using the IMR masses and spins as references
We can now discuss the first hypothesis (i) stated above. Imagine we want to quantify the deviation in each mode’s frequency to put some constraints on an alternative theory. In that case, we have to compare the posterior distributions of frequency and damping time with the QNM values for a BH with the IMR estimated final mass and spin. To simplify, we assume that the parameters estimated from the IMR analysis equal the exact injected values. Results can be seen in Fig. 5, where we subtract the true complex frequency value from each mode. In this figure, we can see that each posterior agrees with the injected value within $2\sigma$. The dashed blue lines mark the quantiles (0.16, 0.84)0.160.84(0.16,\,0.84)( 0.16 , 0.84 ), i.e., the $1\sigma$ distribution.
@@ IV.1.2 Relying only on QNM characterization
Now, to quantify the deviations in this framework, we should translate the differences in mass and spin into deviations in $\omega$ and $\tau$ in terms of $\delta\omega,\delta\tau$, see Eqs. (14), (15). To this end, we assumed that the posterior distribution obtained from the $(2,2,0)$ mode is the “true” description of the remnant BH in GR. We can then compute the QNM spectrum with the derived mass and spin from that mode. This computation is shown in Fig 6, where we observe the posterior of the three complex frequencies $\tilde{\omega}_{lmn}$ computed with the mass and spin derived from the dominant mode $\{\hat{M}_{f220},\hat{a}_{f220}\}$. The GR values are marked with colored crosses on top of the distributions. As stated, deviations might appear in the $(2,2,0)$ mode. Thus, comparing the mass and spin estimated from this BH spectroscopy with those inferred from the full IMR waveforms would be informative.
Given the distributions without deviations, one can measure the one-dimensional deviation for each parameter. To this aim, we need to compare the complex frequency $(lmn)$ posteriors with the values obtained from the mass and spin corresponding to the $(2,2,0)$ mode $(\hat{M}_{f220},\hat{a}_{f220})$ for each mode. This is analogous as comparing Fig. 3 with Fig. 6. The results are shown in the top row of Fig. 7, where we find the distribution of GR complex frequencies computed with the mass and spin derived from the $(2,2,0)$ mode for the $(3,3,0)$ mode in pink, and in green for the $(4,4,0)$ mode. We can easily distinguish them from the non-GR values obtained from the posteriors in blue.
(a) Posterior distribution of $\omega_{330}$ and $\tau_{330}$ against the distribution within GR.
(b) Posterior distribution of $\omega_{440}$ and $\tau_{440}$ against the distribution within GR.
@@
where A and B are two different models, $\mu$ is the estimated mean value and $\sigma$ is the standard deviation. This equation gives the number of standard deviations between two posterior distributions in one dimension. This simple definition can be used as a means to estimate uncertainties in the following. For the injected values of Table 1, the computed standard deviation from GR values is shown in Table 2. Should this be observed, we would have detected a deviation from GR in $\omega_{330}$ with more than 10 standard deviations relative to the $(2,2,0)$ mode. It is also important to note that even though we distinguished a deviation from GR with high precision, the injected value does not correspond to the recovered value, thus revealing a bias.
To summarize, with the hypothesis (ii) this kind of analysis would allow one to differentiate GR from another theory. However, when attempting to constrain an alternative theory, the recovered values of the deviations might lead to misinterpretations. This can be seen in the bottom row of Fig. 7, where the injected value (in black) does not appear consistent with the posterior distribution. For these two bottom figures, we use the blue posterior distributions from the top row and subtract the mean value of the pink and green posteriors respectively. Indeed, the estimated value shown in Fig. 7(c), is inconsistent with $\delta\omega_{330}=0.01$. This is because we used the GR value derived from the mass and spin inferred from the $(2,2,0)$ mode characterization, for which we assumed no deviation. Even if the mass and spin computed from the $(2,2,0)$ mode agree with the true values, the assumption of no deviation in this mode might have strong implications, as any fluctuation on the $(2,2,0)$ mode will translate into fluctuations in the estimated mass and spin and therefore in the subsequent characterization of the $(3,3,0)$ and $(4,4,0)$ modes. Certainly, the computation of the QNMs highly depends on the mass and spin, thus small variations of those intrinsic parameters translate to larger variations on the complex frequency parameter space.
One can avoid this type of discrepancy by using the posterior distribution of the mass and spin inferred from the full IMR instead of the posterior inferred from the $(2,2,0)$ mode. Again, this implies that the IMR analysis should provide unbiased values. In the present analysis, the mass and spin from the $(2,2,0)$ mode were consistent within $2\sigma$ with the injected value.
Figure 8: Posterior distribution for the deviation case. Results without noise injection in red, with noise injection in blue, and injected value with black lines. Overall the distributions agree with the true values, with some fluctuations in the noisy case as expected.
@@
IV.2 Deviation approach
@@
In the following, we discuss the results of the second approach, the deviation template. For this search, we have to define beforehand which QNMs appear in the waveform. We also assume that the dominant mode does not have deviations from GR. Imposing this condition allows us to break the degeneracy between the mode’s fractional deviations from GR and the BH mass and spin. Alternatively, one could fix the mass and the spin, but allow the whole QNM spectrum to present deviations.
@@
In Fig. 8 we show the posterior distribution with and without noise injection in blue and red respectively. Injected values are marked with black lines. We observe the consistency between both results with the true values.
Note that in this approach, the analysis is straightforward. The fractional deviations in the spectrum directly result from the posteriors since the deviations found in each QNM account for the estimated mass and spin by construction. In Fig. 9 we zoom in on the deviations of the $[(3,3,0),(4,4,0)]$ modes and recover the injected values with high accuracy and precision. The uncertainty on the deviations from GR parameters $\delta\omega$ and $\delta\tau$ with this template are listed in Table 3. Note that under the same hypothesis ($ii$) as in previous analysis, i.e. no deviation in the ($2,2,0$) is observed, it is possible to derive constraints on an alternative theory, since the injected values are within the posterior distributions. A caution message is imperative here. The template considered, by construction, does not allow for deviations in the dominant mode. The effect of a fractional deviation in the ($2,2,0$) mode, when not considered in the search template needs further investigation. Nevertheless, in order to constrain an alternative theory the model-independent template might not be enough and specific templates for beyond-GR theories are required.
Table 3: Computed deviation uncertainty from GR for the injected parameters in the deviation approach.
@@
Figure 10: Comparison of the posterior distribution of the fractional deviations for modes (3,3,0) and (4,4,0) for the two different methods. Posterior distributions obtained for the deviation approach are shown in orange and green for the agnostic approach. The black lines intersections mark the injected values.
@@ V Test of GR versus SNR
In what follows, we discuss the SNR needed to claim a deviation from GR with different parameters. To this aim, we will use the deviation template, which provides the best consistency under the assumptions taken. We compute the standard deviation
@@
where $\Gamma_{ii}$ is the Fisher matrix computed as the inner product defined in Eq.(27)
$\Gamma_{ii}=(a(\bm{\theta}_{i})|b(\bm{\theta}_{i})).$ (38)
@@
From Eq. (37) we see how the uncertainty varies as the inverse of the SNR, so naturally the standard deviation in the different parameters decreases as the SNR increases. Consequently, for a given source, it is related to the total mass and the luminosity distance. We can therefore estimate the mass and the redshift needed to claim a deviation from GR with $5\sigma$. Of course, the number of sigmas $N_{\sigma}$ is constrained by the value of the fractional deviation itself, as shown in Eq. (35).
@@
Figure 11: Uncertainty for fractional deviations from GR in $\omega$ and $\tau$ in modes (3,3,0) and (4,4,0) with respect to the source total mass and the redshift. These values are obtained for the fixed parameters listed in Table 1.
@@
We show in Fig. 11 the uncertainty for the parameters with possible deviations from the GR values using the deviation template, such as ($\delta\omega_{330},\delta\tau_{330},\delta\omega_{440},\delta\tau_{440}$). Several assumptions have been made from the beginning of the study, therefore the result we present does not provide a general detection forecast. Nevertheless, this analysis provides a qualitative understanding of LISA’s ability to observe deviations from GR in the ringdown phase of a MBHB coalescence. The uncertainty on the fractional deviations is represented as a function of the source total mass and the redshift which are the dominant contributors to the SNR. We let all other source parameters fixed to the same values listed in Table 1. Consequently, the estimates shown in Fig. 11 are source-dependent, i.e., valid for the particular BH we chose as a case study. Another choice of BH parameters would change this result. Different inclinations, spins, and mass ratios would inevitably change the relative amplitude between QNMs and thus the uncertainty in each mode’s complex frequency.
@@
The color code on the right-hand side of Fig. 11 indicates the value of the uncertainty on the fractional QNM frequencies, obtained with the Fisher matrix for the considered example source. For instance, areas where $\sigma\leq 0.005$ show that LISA should be able to detect deviations from GR in $\delta\omega_{330}$ at the level of 5 standard deviations or more if the departure from GR is of the order of $0.025$ taking $\delta^{GR}=5\sigma$, for sources between $10^{6}$ to $10^{7}$ $M_{\odot}$ through the whole universe, i.e. for any possible redshift. Considering more precision-favorable situations, a deviation from GR of $5\times 10^{-4}$ would be distinguishable for sources below redshift 1 and total mass of the order of $10^{7}\,M_{\odot}$. At first glance, one could conclude that the most severe limits will come from the deviations or lack of them in the frequencies $\delta\omega_{lmn}$ because of the high sensitivity of LISA to frequency variations. The expected population of MBHB for heavy seeds encloses sources in the range $[10^{4}-10^{7}]\,M_{\odot}$ up to redshift 10 approximately. This range is extended to lower-mass sources in the case of light seeds [43, 78, 79]. Hence, even if LISA cannot observe some of these golden sources, the expectation to test the no-hair theorem and GR looks very promising for “nearby” sources ranging in the mass interval $[10^{6}-10^{7}]\,M_{\odot}$.

### v2

@@ III.2 Templates
We consider two approaches where the recovery template takes different forms: the agnostic approach and the deviations approach. In the agnostic approach, we assume that the ringdown waveform is described by
@@
Here, complex frequencies, amplitudes, and phases can take any value. Note also that any dependence on the spheroidal harmonic is absorbed in the amplitude and the phase. In this way, no mode mixing is specified. For example, in this approach, it would not be possible to know how much of the $(2,2,0)$ QNM contribution comes from the spherical harmonic $(2,2)$ or $(3,2)$. It differs from Eq. (1) as no assumption is made on the value of the complex frequency nor the spherical contribution to any QNM. Despite the fixed number of modes $k$, we can call this approach “agnostic”.
In the deviations approach, we assume the framework of GR as baseline but allow for a small “deviation” in the complex frequencies
@@ III.3 Bayesian analysis
where $p()$ are the probability density functions, $\bm{\theta}$ is the vector of the source physical parameters, $M$ is the model or any other feature considered. In the numerator, we have the likelihood $\mathcal{L}(\theta)=p(d|\bm{\theta},M)$ and the prior of the parameters $\pi(\theta)=p(\bm{\theta}|M)$. In the numerator, $\mathcal{Z}=p(d|M)$ is the evidence, which is computed as the integral of the likelihood over the whole parameter’s hyper-volume. For a noise with a covariance matrix $\mathbf{C}$, the likelihood takes the form
@@
for a stationary noise. We drop the dependence on the model $M$ and we use the definition of the inner product in the time domain
@@
Table 2: Priors on the parameters estimated for both approaches
Parameter Priors for agnostic Priors for deviations
$A_{lmn}$ log uniform [-23,-16] log uniform [-23,-16]
$\phi_{lmn}$ uniform [0,$2\pi$] uniform [0,$2\pi$]
$\omega^{\prime}_{lmn}$ uniform [$10^{-5}$,0.1] -
$\tau^{\prime}_{lmn}$ uniform [1,$10^{5}$] -
$\delta\omega_{lmn}$ [-0.2, 0.2]
$\delta\tau_{lmn}$ [-0.2, 0.2]
$M_{f}$ [0.9, 1.1] ×$\times M_{f}$
$a_{f}$ [0.9, 1.1] ×$\times a_{f}$
@@
where $\rho(k)$, $k=|i-j|$, is the autocovariance function (ACF) that can be estimated from noise-only data in the time domain with a length longer than $N$ or as the inverse Fourier transform of the PSD. In our case, we use the latter option, generating the ACF from the LISA Science Requirements Document (SCiRD) [84] PSD
@@
Then, the inner product of Eq. (27) becomes a fast product of vectors
@@
Figure 2: Posterior distribution of the pairs of complex frequencies in the spectrum map. Each mode ‘k’ is associated with one color, purple, pink, or green. The spectrum of the true BH is represented with colored crosses with their QNM label nearby.
@@ IV.1 Agnostic approach
[…]
@@
In Fig. 1, we present the posterior distribution of the injection without noise in red, with noise in blue, and the injected values with black lines. Remember that the values of the deviations injected in the waveform have been introduced in Table 1. We show 12 parameters, 4 for each QNM. In general, the Gaussian distributions converge to the true values, with some minor fluctuations in the noisy case, as expected. In the following, we will discuss the results from the noisy dataset only. In the figure we have already labeled the name of the QNMs “$k$” since we know them from the injection. However, we will not know which modes are present in the future LISA data analysis. Therefore, one must first find the QNM corresponding to each label $k=lmn$.
@@
Figure 4: Computed mass for spins in the range [0.8-0.9] with Eq. (34) with a $99\%$ confidence level, from the estimated mean values of ($\omega_{lmn},\tau_{lmn}$)
One possible way to check for consistency between QNMs is to trace back the mass and spin of the remnant BH, using for example the parametrization from [27]:
@@ IV.1.1 Using the IMR masses and spins as references
We can now discuss the first hypothesis (i) stated above. Imagine we want to quantify the deviation in each mode’s frequency to put some constraints on an alternative theory. In that case, we have to compare the posterior distributions of frequency and damping time with the QNM values for a BH with respect to the IMR estimated final mass and spin. To simplify, we assume that the parameters estimated from the IMR analysis equal the exact injected values. Results can be seen in Fig. 5, where we subtract the true complex frequency value from each mode. In this figure, we can see that each posterior agrees with the injected value within $2\sigma$. The dashed blue lines mark the quantiles (0.16, 0.84)0.160.84(0.16,\,0.84)( 0.16 , 0.84 ), i.e., the $1\sigma$ distribution.
@@ IV.1.2 Relying only on QNM characterization
Now, to quantify the deviations in this framework, we should translate the differences in mass and spin into deviations in $\omega$ and $\tau$ in terms of $\delta\omega,\delta\tau$, see Eqs. (14), (15). To this end, we assumed that the posterior distribution obtained from the $(2,2,0)$ mode is the “true” description of the remnant BH in GR. We can then compute the QNM spectrum with the derived mass and spin from that mode. This computation is shown in Fig 6, where we observe the posterior of the three complex frequencies $\tilde{\omega}_{lmn}$ computed with the mass and spin derived from the dominant mode $\{\hat{M}_{f220},\hat{a}_{f220}\}$. The GR values are marked with colored crosses on top of the distributions.
Given the distributions without deviations, one can measure the one-dimensional deviation for each parameter. To this aim, we need to compare the complex frequency $(\tilde{\omega}_{lmn})$ posteriors with the QNM values obtained from the mass and spin corresponding to the $(2,2,0)$ mode $(\hat{M}_{f220},\hat{a}_{f220})$ for each mode. This is analogous as comparing Fig. 3 with Fig. 6. The results are shown in the top row of Fig. 7, where we find the distribution of GR complex frequencies computed with the mass and spin derived from the $(2,2,0)$ mode for the $(3,3,0)$ mode in pink, and in green for the $(4,4,0)$ mode. We can easily distinguish them from the non-GR values obtained from the sampler in blue.
@@
where A and B are two different models, $\mu$ is the estimated mean value and $\sigma$ is the standard deviation. This equation gives the number of standard deviations between two posterior distributions in one dimension. This simple definition can be used as a means to estimate uncertainties in the following. For the injected values of Table 1, the computed standard deviation from GR values is shown in Table 3. Should this be observed, we would have detected a deviation from GR in $\omega_{330}$ with more than 10 standard deviations relative to the $(2,2,0)$ mode. It is also important to note that even though we distinguished a deviation from GR with high precision, the injected GR value does not correspond to the recovered GR value, thus revealing a bias.
@@
To compute the values of the fractional deviations for each mode, one can use the blue posterior distributions from the top row and subtract the mean value of the pink and green posteriors respectively. Performing such computation leads to the figures in the bottom row in Fig. 7. As expected from the previous remark on the GR inconsistent values, the injected parameter does not perfectly agree with the posterior distribution. Indeed, the estimated value shown in Fig. 7(c), is inconsistent with $\delta\omega_{330}=0.01$. This is because we used the GR value derived from the mass and spin inferred from the $(2,2,0)$ mode characterization, for which we assumed no deviation. Even if the mass and spin computed from the $(2,2,0)$ mode agree with the true values, the assumption of no deviation in this mode might have strong implications, as any fluctuation on the $(2,2,0)$ mode will translate into fluctuations in the estimated mass and spin and therefore in the subsequent characterization of the $(3,3,0)$ and $(4,4,0)$ modes. Certainly, the computation of the QNMs highly depends on the mass and spin, thus small variations of those intrinsic parameters translate to larger variations on the complex frequency parameter space. To summarise, with the hypothesis (ii), this kind of analysis would allow one to differentiate GR from another theory. However, when attempting to constrain an alternative theory, the recovered values of the deviations might lead to misinterpretations.
One can avoid this type of discrepancy by using the posterior distribution of the mass and spin inferred from the full IMR instead of the posterior inferred from the $(2,2,0)$ mode. Again, this implies that the IMR analysis should provide unbiased values. Keep in mind that, in the present analysis, the mass and spin from the $(2,2,0)$ mode were consistent within $2\sigma$ with the injected values of final mass and final spin. Nevertheless, deviations might appear in the $(2,2,0)$ mode. Thus, comparing the mass and spin estimated from this BH spectroscopy with those inferred from the full IMR waveforms would be informative.
Figure 8: Posterior distribution for the deviations case. Results without noise injection in red, with noise injection in blue, and injected value with black lines. Overall the distributions agree with the true values, with some fluctuations in the noisy case as expected.
IV.2 Deviations approach
@@
In the following, we discuss the results of the second approach, the deviations template. For this search, we have to define beforehand which QNMs appear in the waveform. We also assume that the dominant mode does not have deviations from GR. Imposing this condition allows us to break the degeneracy between the mode’s fractional deviations from GR and the BH mass and spin. Alternatively, one could fix the mass and the spin, and allow the whole QNM spectrum to present deviations.
@@
In Fig. 8 we show the posterior distribution with and without noise injection in blue and red respectively. Injected values are marked with black lines. We observe the consistency between both results with the true values. In the following, we analyze the results from the dataset with noise injection.
Note that in this approach, the analysis is straightforward. The fractional deviations in the spectrum directly result from the posteriors since the deviations found in each QNM account for the estimated mass and spin by construction. In Fig. 9 we zoom in on the deviations of the $[(3,3,0),(4,4,0)]$ modes and recover the injected values with high accuracy and precision. The uncertainty on the deviations from GR parameters $\delta\omega$ and $\delta\tau$ with this template are listed in Table 4. Note that under the same hypothesis ($ii$) as in the previous analysis, i.e. no deviation in the ($2,2,0$) is observed, it is possible to derive constraints on an alternative theory, since the injected values are within the posterior distributions. A caution message is imperative here. The template considered, by construction, does not allow for deviations in the dominant mode. The effect of a fractional deviation in the ($2,2,0$) mode, when not considered in the search template needs further investigation. Nevertheless, to constrain an alternative theory the model-independent template might not be enough, and specific templates for beyond-GR theories are required.
Table 4: Computed deviation uncertainty from GR for the injected parameters in the deviations approach.
@@
Figure 10: Comparison of the posterior distribution of the fractional deviations for modes (3,3,0) and (4,4,0) for the two different methods. Posterior distributions obtained for the deviations approach are shown in orange and green for the agnostic approach. The black lines intersections mark the injected values.
@@ V Test of GR versus SNR
In what follows, we discuss the uncertainty in the fractional deviations for a source with different SNR across the LISA band. To this aim, we will use the deviations template, which provides the best consistency under the assumptions taken. We compute the standard deviation
@@
where $\Gamma_{ii}$ is the Fisher matrix computed as the inner product defined in Eq.(27) of the partial derivatives with respect to the parameter $\bm{\theta}_{i}$
$\Gamma_{ii}=\left(\frac{\partial a(\bm{\theta})}{\partial\bm{\theta}_{i}}\Big{|}\frac{\partial b(\bm{\theta})}{\partial\bm{\theta}_{i}}\right).$ (38)
@@
From Eq. (37) we see how the uncertainty varies as the inverse of the SNR, so naturally the standard deviation in the different parameters decreases as the SNR increases. Consequently, for a given source, the uncertainty is related to the total mass and the luminosity distance. To indistinctly translate luminosity distance ($D_{l}$) to redshift (z), we assume cosmological parameters widely used within the LISA Data Challenge (LDC) group, i.e. flat $\Lambda_{CDM}$, with Hubble constant $H_{0}=67.74$ and matter density $\Omega_{M}=0.3075$, in agreement with Planck’s results [94, 95]. We can therefore estimate the mass and the redshift needed to claim a deviation from GR with $5\sigma$. Of course, the number of sigmas $N_{\sigma}$ is constrained by the value of the fractional deviation itself, as shown in Eq. (35).
@@
Figure 11: Uncertainty for fractional deviations from GR in $\omega$ and $\tau$ in modes (3,3,0) and (4,4,0) with respect to the source total mass and the redshift. These values are obtained for the fixed parameters listed in Table 1. The dotted black lines correspond to contours for a ringdown-only SNR equal to $[8,20,50,100,250,500,1000,3000]$.
@@
We show in Fig. 11 the uncertainty for the parameters with possible deviations from the GR values, such as ($\delta\omega_{330},\delta\tau_{330},\delta\omega_{440},\delta\tau_{440}$), using the deviations template. Several assumptions have been made from the beginning of the study, therefore the result we present does not provide a general detection forecast. Nevertheless, this analysis provides a qualitative understanding of LISA’s ability to observe deviations from GR in the ringdown phase of a MBHB coalescence. The uncertainty on the fractional deviations is represented as a function of the source total mass and the redshift which are the dominant contributors to the SNR. We let all other source parameters fixed to the same values listed in Table 1. Consequently, the estimates shown in Fig. 11 are source-dependent, i.e., valid for the particular BH we chose as a case study. Another choice of BH parameters would change this result. Different inclinations, spins, and mass ratios would inevitably change the relative amplitude between QNMs and thus the uncertainty in each mode’s complex frequency. For comparison, the value of the ringdown’s SNR is also shown in the figure as contour dotted lines for the values $[8,20,50,100,250,500,1000,3000]$.
@@
The color code on the right-hand side of Fig. 11 indicates the value of the uncertainty on the fractional QNM frequencies, obtained with the Fisher matrix for the considered example source. For instance, areas where $\sigma\leq 0.005$ show that LISA should be able to detect deviations from GR in $\delta\omega_{330}$ at the level of 5 standard deviations or more if the departure from GR is of the order of $0.025$ taking $\delta^{GR}=5\sigma$, for sources between $10^{6}$ to $10^{7}$ $M_{\odot}$ through the whole universe, i.e. for any possible redshift. Considering more precision-favorable situations, a deviation from GR of $5\times 10^{-5}$ would be distinguishable for sources below redshift 1 and total mass of the order of $10^{7}\,M_{\odot}$. At first glance, one could conclude that the most severe limits will come from the deviations or lack of them in the frequencies $\delta\omega_{lmn}$ because of the high sensitivity of LISA to frequency variations. The expected population of MBHB for heavy seeds encloses sources in the range $[10^{4}-10^{7}]\,M_{\odot}$ up to redshift 10 approximately. This range is extended to lower-mass sources in the case of light seeds [55, 96, 97]. Hence, even if LISA cannot observe some of these golden sources, the expectation to test the no-hair theorem and GR looks very promising for “nearby” sources ranging in the mass interval $[10^{6}-10^{7}]\,M_{\odot}$.
[…]
Table 6: Uncertainty computed with the Fisher matrix (second column) and obtained with the sampler (third column) for a low SNR ($\sim 15$) source.
$\sigma_{\mathrm{FM}}$ $\sigma_{\mathrm{sampler}}$
$\delta\omega_{330}$ $0.036$ $0.044$
$\delta\tau_{330}$ $0.465$ $0.127$
$\delta\omega_{440}$ $0.243$ $0.066$
$\delta\tau_{440}$ $1.881$ $0.156$